# Efficient Dynamics Modeling in Interactive Environments with Koopman Theory

### v1

@@ Appendix C Proposed modified dynamics model in RL and Planning
C.1 Koopman TDMPC
@@
where $R_{\theta}$, $\pi_{\theta}$, $Q_{\theta}$ are respectively reward, policy and Q prediction networks, where $y_{t+k}=r_{t+k}+Q_{\theta^{-}}(\hat{x}_{t+k+1},\pi_{\theta}(\hat{x}_{t+k}))$ is the 1-step bootstrapped TD target. Moreover, the policy network is learned from the latent representation by maximizing the Q value at each time step. This additional policy network helps to provide the next action for the TD target and a heuristic for planning. For details on the planning algorithm and its implementation, see [8].
C.2 Koopman Self-Predictive Representations
@@
Here, $a_{t+1}$ is sampled from the policy $\pi$. The policy is learned from the representations using Soft Actor-Critic [SAC; 51]. As opposed to SPR, we do not use moving averages for the target encoder parameters and simply stop gradients through the target encoders and denote it as $e_{\theta^{-}}$. Moreover, we drop the consistency term in the Koopman space in Eq. 15 as we empirically observed adding consistency both in Koopman observable ($x$), and Q-learning space ($z$) promotes collapse and makes training unstable, resulting in a higher variance in the model’s performance.
Appendix D Ablation study on Initialization
@@
We now perform an ablation study on the different possible initialization methods presented in Section 3.2.2. We report the results for halfcheetah-expert-v2, walker-2d-expert-v2 and hopper-expert-v2, in Table 5, 5 and 5 respectively. We see a constant value of $-0.2$ for $\mu_{i}$, and increasing the order of frequency for $\omega_{i}$ gives a consistent performance boost over different environments.
@@
Appendix E Ablation on the dimensionality of Linear latent space
@@
Appendix F More details on dynamics models
For all the experiments, we used a model with around $\sim 500k$ trainable parameters to ensure that the improved compute efficiency during training is not due to the model having fewer parameters. To incorporate the causal transformer architecture with that many parameters, we reduced the embedding dimension of both the state and action embedding to $64$. We further reduced the number of heads to $4$ to design a causal Transformer model that matches the size of our Koopman model. For the MLP-based model, we keep these dimensions the same as the Koopman one. Moreover, all our experiments in Table 2 use one Nvidia A100 GPU for training.
Appendix G Koopman dynamics model in jax
We provide the Jax code snippet to efficiently implement our Koopman-based dynamics model.
@@ Appendix G Koopman dynamics model in jax
import jax
import jax.numpy as jnp
from jax.nn.initializers import normal
from flax import linen as nn
# Main class to implement Diagonal Koopman Dynamics model
class BatchedDiagonalKoopmanDynamicsModel(nn.Module):
    state_dim: int
    action_emb_dim: int
    real_init_type: str = ’constant’
    real_init_value: float = -0.5
    im_init_type: str = ’increasing_freq’
    activations: Callable[[jnp.ndarray], jnp.ndarray] = nn.relu
    def init_koopman_imaginary_params(self):
        if self.im_init_type == ’increasing_freq’:
            self.K_im = self.param(
            "K_im", increasing_im_init(), (self.state_dim // 4,)
        elif self.im_init_type == ’random’:
            self.K_im = self.param(
            "K_im", random_im_init(), (self.state_dim // 4,)
        else:
            raise ValueError
    def init_koopman_real_params(self):
        if self.real_init_type == ’constant’:
            self.K_real = self.real_init_value
        elif self.real_init_type == ’learnable’:
            self.K_r = self.param(
            "K_real", nn.initializers.ones, (self.state_dim // 4,)
            ) * self.real_init_value
            self.K_real = jnp.clip(self.K_r, -0.4, -0.1)
        else:
            raise ValueError
    def setup(self):
        # model parameters
        self.init_koopman_real_params()
        self.init_koopman_imaginary_params()
        self.action_encoder = MLP(
        [128, self.action_emb_dim * 2],
        activations=self.activations
        self.L = self.param(
        "L", normal(0.1), (self.action_emb_dim, self.state_dim // 2, 2)
        # Step parameter
        self.step = jnp.exp(self.param("log_step", log_step_init(), (1,)))
        self.K_complex = jnp.concatenate(
            [self.K_real + 1j * self.K_im,
            self.K_real - 1j * self.K_im],
            axis=-1
        self.L_complex = self.L[:, :, 0] + 1j * self.L[:, :, 1]
        self.K_dis, self.L_dis = discretize(
            self.K_complex, self.L_complex, self.step
    def __call__(self, actions, start_state_rep):
        if len(actions.shape) == 2:
            return koopman_forward_single(
                self.K_dis, self.L_dis,
                self.action_encoder(actions), start_state_rep
        else:
            Vand_K = jnp.vander(
                self.K_dis, actions.shape[1] + 1, increasing=True
            return koopman_forward(
                Vand_K, self.L_dis,
                self.action_encoder(actions), start_state_rep
@@
def causal_convolution(u, K):
    assert K.shape[0] == u.shape[0]
    ud = jnp.fft.fft(jnp.pad(u, (0, K.shape[0])))
    Kd = jnp.fft.fft(jnp.pad(K, (0, u.shape[0])))
    out = ud * Kd
    return jnp.fft.ifft(out)[: u.shape[0]]
def compute_measurement_block(Ker, action_emb, init_state):
    return causal_convolution(action_emb, Ker[:-1]) + Ker[1:] * init_state
def compute_measurement(Ker, action_emb, init_state):
    return jax.vmap(
            compute_measurement_block,
            in_axes=(0, 1, 0),
            out_axes=1
        )(Ker, action_emb, init_state)
def discretize(K, L, step):
    return jnp.exp(step*K), (jnp.exp(step*K)-1)/K * L
def log_step_init(dt_min=0.001, dt_max=0.1):
    def init(key, shape):
        return jax.random.uniform(key, shape) * (
            jnp.log(dt_max) - jnp.log(dt_min)
        ) + jnp.log(dt_min)
    return init
def increasing_im_init():
    def init(key, shape):
        return jnp.pi * jnp.arange(shape[0])
    return init
def random_im_init():
    def init(key, shape):
        return jax.random.uniform(key, shape)
    return init
@jit
def koopman_forward(Vand_K, L, actions_emb, start_state_rep):
    action_emb_dim = actions_emb.shape[-1]
    state_dim = start_state_rep.shape[-1]
    actions_emb_complex = actions_emb[:, :, :action_emb_dim//2] \
                          + 1j * actions_emb[:, :, action_emb_dim//2:]
    start_state_rep_complex = start_state_rep[:, :state_dim // 2] \
                              + 1j * start_state_rep[:, state_dim // 2:]
    predicted_state_rep_complex = jax.vmap(
        compute_measurement,
        in_axes=(None, 0, 0),
        out_axes=0
    )(Vand_K, actions_emb_complex @ L, start_state_rep_complex)
    predicted_state_rep = jnp.concatenate([
        jnp.real(predicted_state_rep_complex),
        jnp.imag(predicted_state_rep_complex)
    ], -1)
    return predicted_state_rep
@jit
def koopman_forward_single(K, L, action_emb, start_state_rep):
    action_emb_dim = action_emb.shape[-1]
    state_dim = start_state_rep.shape[-1]
    action_emb_complex = action_emb[:, :action_emb_dim//2] \
                         + 1j * action_emb[:, action_emb_dim//2:]
    start_state_rep_complex = start_state_rep[:, :state_dim // 2] \
                              + 1j * start_state_rep[:, state_dim // 2:]
    predicted_state_rep_complex = K[None, :] * start_state_rep_complex + \
                                 .action_emb_complex @ L
    predicted_state_rep = jnp.concatenate([
        jnp.real(predicted_state_rep_complex),
        jnp.imag(predicted_state_rep_complex)
    ], -1)
    return predicted_state_rep

### v4

@@ Appendix C Proposed modified dynamics model in RL and Planning
C.1 Reinforcement Learning
Value-based RL algorithms have two steps of policy evaluation and policy improvement. For example, $Q$-learning estimates the optimal action-value function, denoted by $Q^{*}(s,a)$, that represents the maximum expected cumulative reward achievable by taking action $a$ in state $s$ and improves the policy by setting it to $arg\max_{a}Q^{*}(s,a)$ (Mnih et al., 2013; 2015). A policy-based RL algorithm directly optimizes the policy to maximize the expected return. Policy gradient methods update the policy by adjusting the policy function’s parameters by estimating the gradient of $J$. Actor-critic methods combine Q-learning and policy gradient ones, such that the Q-network (critic) learns the action value under the current policy, and the policy network (actor) tries to maximize it Haarnoja et al. (2018), Lillicrap et al. (2016), Schulman et al. (2017).
C.2 Forward dynamics modeling in RL
Dynamics models have been instrumental in attaining impressive results on various tasks such as Atari (Schrittwieser et al., 2020, Hafner et al., 2020) and continuous control (Hafner et al., 2020, Jiang et al., 2020, Sikchi et al., 2021, Lowrey et al., 2019). By building a model of environment dynamics, model-based RL can generate trajectories used for training the RL algorithm. This can significantly reduce the sample complexity in comparison to model-free techniques. However, in model-based RL, inaccurate long-term predictions can generate poor trajectory rollouts and lead to incorrect expected return calculations, resulting in misleading policy updates. Forward dynamics modeling has also been successfully applied to model-free RL to improve the sample efficiency of the existing model-free algorithms. These methods use it to design self-supervised auxiliary losses for representation learning using consistency in forward dynamics in the latent and the observation space (Jaderberg et al., 2017, Schwarzer et al., 2020, Srinivas et al., 2020).
C.3 Model-based Planning
where $\gamma$ is typically set to 1, that is, there is no discounting. Heuristic population-based methods, such as a cross entropy method Rubinstein & Kroese (2004) are often used for dynamic planning. These methods perform a local trajectory optimization problem, corresponding to the optimization of a cost function up to a certain number of time steps in the future (Williams et al., 2015; 2018, Okada & Taniguchi, 2020). In contrast to Q-learning, this is myopic and can only plan up to a certain horizon, which is predetermined by the algorithm.
One can also combine MPC with RL to approximate long-term returns of trajectories that can be calculated by bootstrapping the value function of the terminal state. In particular, methods like TD-MPC (Hansen et al., 2022) and LOOP (Silkchi2020LOOP) combine value learning with planning using MPC. The learned value function is used to bootstrap the trajectories that are used for planning using MPC. Additionally, they learn a policy using the Deep Deterministic Policy Gradient (Lillicrap et al., 2016) or Soft Actor-Critic (SAC) (Haarnoja et al., 2018) to augment the planning algorithm with good proposal trajectories. Alternative search method such as Monte Carlo Tree search (Coulom, 2006) is used for planning in discrete action spaces. The performance of the model-based planning method heavily relies on the accuracy of the learned model.
C.4 Koopman TDMPC
@@
where $R_{\theta}$, $\pi_{\theta}$, $Q_{\theta}$ are respectively reward, policy and Q prediction networks, where $y_{t+k}=r_{t+k}+Q_{\theta^{-}}(\hat{x}_{t+k+1},\pi_{\theta}(\hat{x}_{t+k}))$ is the 1-step bootstrapped TD target. Moreover, the policy network is learned from the latent representation by maximizing the Q value at each time step. This additional policy network helps to provide the next action for the TD target and a heuristic for planning. For details on the planning algorithm and its implementation, see Hansen et al. (2022).
C.5 Koopman Self-Predictive Representations
@@
Here, $a_{t+1}$ is sampled from the policy $\pi$. The policy is learned from the representations using Soft Actor-Critic (SAC; Haarnoja et al., 2018). As opposed to SPR, we do not use moving averages for the target encoder parameters and simply stop gradients through the target encoders and denote it as $e_{\theta^{-}}$. Moreover, we drop the consistency term in the Koopman space in Eq. 10 as we empirically observed adding consistency both in Koopman observable ($x$), and Q-learning space ($z$) promotes collapse and makes training unstable, resulting in a higher variance in the model’s performance.
Appendix D Comparison with DSSM
While our proposed Koopman model may look similar to a DSSM Gu et al. (2022a), Gupta et al. (2022), Mehta et al. (2022), there are four major differences. First, our model is specifically derived for dynamics modeling with control input using Koopman theory and not for sequence or time series modeling. Our motivation is to make the dynamics and gradients through time stable, whereas, for DSSMs, it is to approximate the 1D convolution kernels derived from HiPPO Gu et al. (2020) theory. Second, a DSSM gives a way to design a cell that is combined with non-linearity and layered to get a non-linear sequence model. In contrast, our Koopman-based model shows that simple linear latent dynamics can be sufficient to model complex non-linear dynamical systems with control. Third, DSSMs never explicitly calculate the latent states and even ignore the starting state. Our model works in the state space, where the starting state is crucial to backpropagate gradients through the state encoder. Fourth, a DSSM learns structured convolution kernels for 1D to 1D signal mapping so that for higher dimensional input, it has multiple latent dynamics models running under the hood, which are implemented as convolutions. In contrast to this, our model runs a single linear dynamics model for any dimensional input.
Appendix E Detailed results from the main text
In this section, we provide the numerical values in Table 1 and Table 2 of the results used in Fig. 3. Table 3 provides the numerical values for training speed used in Fig. 4. In Fig. 7 and Fig. 8, we further provide the environment-wise returns evaluated after every $5000$ training steps for the planning and RL experiments used in Fig. 5 and Fig. 6.
Table 1: Forward state prediction error in Offline Reinforcement Learning environments. The reported number is the mean square error of the state prediction over a horizon of 100 environment steps. The results include standard deviation over 3 runs.
Offline Datasets State Prediction Error ($\downarrow$)
MLP Transformer GRU DSSM Koopman
hopper-expert-v2 0.500 $\pm$ 0.080 0.060 $\pm$ 0.007 0.006 $\pm$ 0.001 0.030 $\pm$ 0.008 0.007 $\pm$ 0.000
hopper-medium-expert-v2 0.550 $\pm$ 0.120 0.063 $\pm$ 0.009 0.011 $\pm$ 0.004 0.050 $\pm$ 0.008 0.012 $\pm$ 0.004
hopper-full-replay-v2 0.880 $\pm$ 0.150 0.240 $\pm$ 0.070 0.050 $\pm$ 0.005 0.180 $\pm$ 0.050 0.070 $\pm$ 0.004
walker2d-expert-v2 0.300 $\pm$ 0.060 0.130 $\pm$ 0.060 0.100 $\pm$ 0.005 0.140 $\pm$ 0.020 0.090 $\pm$ 0.008
walker2d-medium-expert-v2 1.320 $\pm$ 0.210 0.310 $\pm$ 0.070 0.190 $\pm$ 0.050 0.320 $\pm$ 0.070 0.220 $\pm$ 0.040
walker2d-full-replay-v2 4.500 $\pm$ 2.100 3.300 $\pm$ 0.350 0.850 $\pm$ 0.080 1.200 $\pm$ 0.090 0.910 $\pm$ 0.050
halfcheetah-expert-v2 2.860 $\pm$ 1.100 0.840 $\pm$ 0.160 0.280 $\pm$ 0.050 0.520 $\pm$ 0.050 0.300 $\pm$ 0.060
halfcheetah-medium-expert-v2 0.590 $\pm$ 0.090 0.540 $\pm$ 0.071 0.501 $\pm$ 0.031 0.910 $\pm$ 0.040 0.568 $\pm$ 0.040
halfcheetah-full-replay-v2 0.850 $\pm$ 0.200 0.930 $\pm$ 0.140 0.750 $\pm$ 0.030 0.950 $\pm$ 0.060 0.700 $\pm$ 0.040
Table 2: Forward reward prediction error in Offline Reinforcement Learning environments. The reported number is the mean square error of the reward prediction over a horizon of 100 environment steps. The results include standard deviation over 3 runs.
Offline Datasets Reward Prediction Error ($\downarrow$)
MLP Transformer GRU DSSM Koopman
hopper-expert-v2 0.080 $\pm$ 0.005 0.030 $\pm$ 0.002 0.001 $\pm$ 0.000 0.002 $\pm$ 0.000 0.001 $\pm$ 0.000
hopper-medium-expert-v2 0.260 $\pm$ 0.050 0.046 $\pm$ 0.008 0.001 $\pm$ 0.000 0.005 $\pm$ 0.000 0.002 $\pm$ 0.000
hopper-full-replay-v2 0.400 $\pm$ 0.100 0.110 $\pm$ 0.040 0.009 $\pm$ 0.005 0.020 $\pm$ 0.004 0.012 $\pm$ 0.003
walker2d-expert-v2 0.015 $\pm$ 0.003 0.004 $\pm$ 0.001 0.004 $\pm$ 0.001 0.006 $\pm$ 0.001 0.003 $\pm$ 0.000
walker2d-medium-expert-v2 0.190 $\pm$ 0.030 0.032 $\pm$ 0.009 0.009 $\pm$ 0.003 0.015 $\pm$ 0.003 0.011 $\pm$ 0.002
walker2d-full-replay-v2 0.520 $\pm$ 0.080 0.220 $\pm$ 0.020 0.080 $\pm$ 0.015 0.091 $\pm$ 0.005 0.090 $\pm$ 0.008
halfcheetah-expert-v2 0.160 $\pm$ 0.050 0.070 $\pm$ 0.005 0.045 $\pm$ 0.004 0.120 $\pm$ 0.020 0.040 $\pm$ 0.006
halfcheetah-medium-expert-v2 0.230 $\pm$ 0.050 0.120 $\pm$ 0.070 0.050 $\pm$ 0.007 0.150 $\pm$ 0.007 0.060 $\pm$ 0.005
halfcheetah-full-replay-v2 0.170 $\pm$ 0.045 0.110 $\pm$ 0.052 0.105 $\pm$ 0.009 0.205 $\pm$ 0.011 0.091 $\pm$ 0.006
Dynamics Model Type Training sequence length 10 50 100 200 500 MLP 478.00 115.54 55.85 28.02 11.20 Transformer 418.00 120.20 73.34 34.48 14.92 GRU 420.00 101.20 42.02 17.40 8.91 DSSM 530.32 148.50 76.10 37.12 17.20 Koopman 565.43 160.76 86.49 59.65 24.56
Figure 7: Comparison of our Koopman-based dynamics model (with a horizon of 20) with an MLP-based dynamics model of vanilla TD-MPC Hansen et al. (2022). The results are shown with 5 random seeds with the standard deviation shown using the shaded regions.
Figure 8: Comparison of vanilla RL (SAC Haarnoja et al. (2018)) and its integration with an MLP-based (SPR Schwarzer et al. (2020)) and a Koopman-based dynamics model for incorporating self-predictive representations in the DeepMind Control Suite. Shaded regions represent the standard deviation.
Appendix F More details on the baseline dynamics models
For all the experiments, we used a model with around $\sim 500k$ trainable parameters to ensure that the improved compute efficiency during training is not due to the model having fewer parameters. To incorporate the causal transformer architecture with that many parameters, we reduced the embedding dimension of both the state and action embedding to $64$. We further reduced the number of heads to $4$ to design a causal Transformer model that matches the size of our Koopman model. For the GRU and MLP-based model, we keep these dimensions the same as the Koopman one. We use a hidden dimension size of 64 for the GRU based-model in order to keep the parameters similar. Moreover, all our experiments in Table 3 use one Nvidia A100 GPU for training.
Appendix G Ablation study on Initialization
@@
We now perform an ablation study on the different possible initialization methods presented in Section 3.2.2. We report the results for halfcheetah-expert-v2, walker-2d-expert-v2 and hopper-expert-v2, in Table 4, 5 and 6 respectively. We see a constant value of $-0.2$ for $\mu_{i}$, and increasing the order of frequency for $\omega_{i}$ gives a consistent performance boost over different environments.
@@
Appendix H Ablation study on horizon for Planning and Model-free RL
We investigate the role of horizon (or rollout length) in both model-based planning (Section 4.2.1) and model-free RL (Section 4.2.2) for our Koopman-based dynamics model. The corresponding results are available in Fig. 9 and Fig. 10. We observe that for model-based planning, the performance improves with longer horizon. However, for model-free RL, the performance is higher with a prediction horizon of 5 and deteriorates for 20. This phenomenon is consistent with the results reported in Figure 4 of Schwarzer et al. (2020).
Figure 9: Ablations of our Koopman-based dynamics model for horizon lengths of 5 and 20. The results are over 5 random seeds for each environment. Higher Mean &\&& IQM and lower Optimality Gap is better.
Figure 10: Ablations of our Koopman-based dynamics model for horizon lengths of 5 and 20 for incorporating self-predictive representations with vanilla SAC Haarnoja et al. (2018) in the DeepMind Control Suite. The results are over 5 random seeds for each environment. Higher Mean &\&& IQM and lower Optimality Gap is better.
Appendix I Ablation on the dimensionality of Linear latent space
@@
Appendix J Koopman dynamics model in jax
We provide the Jax code snippet to efficiently implement our Koopman-based dynamics model. Our code can be found in https://github.com/arnab39/koopman-dynamics.
@@
import jax import jax.numpy as jnp from jax.nn.initializers import normal from flax import linen as nn # Main class to implement Diagonal Koopman Dynamics model class BatchedDiagonalKoopmanDynamicsModel(nn.Module): state_dim: int action_emb_dim: int real_init_type: str = ’constant’ real_init_value: float = -0.5 im_init_type: str = ’increasing_freq’ activations: Callable[[jnp.ndarray], jnp.ndarray] = nn.relu def init_koopman_imaginary_params(self): if self.im_init_type == ’increasing_freq’: self.K_im = self.param( "K_im", increasing_im_init(), (self.state_dim // 4,) ) elif self.im_init_type == ’random’: self.K_im = self.param( "K_im", random_im_init(), (self.state_dim // 4,) ) else: raise ValueError def init_koopman_real_params(self): if self.real_init_type == ’constant’: self.K_real = self.real_init_value elif self.real_init_type == ’learnable’: self.K_r = self.param( "K_real", nn.initializers.ones, (self.state_dim // 4,) ) * self.real_init_value self.K_real = jnp.clip(self.K_r, -0.4, -0.1) else: raise ValueError def setup(self): # model parameters self.init_koopman_real_params() self.init_koopman_imaginary_params() self.action_encoder = MLP( [128, self.action_emb_dim * 2], activations=self.activations ) self.L = self.param( "L", normal(0.1), (self.action_emb_dim, self.state_dim // 2, 2) ) # Step parameter self.step = jnp.exp(self.param("log_step", log_step_init(), (1,))) self.K_complex = jnp.concatenate( [self.K_real + 1j * self.K_im, self.K_real - 1j * self.K_im], axis=-1 ) self.L_complex = self.L[:, :, 0] + 1j * self.L[:, :, 1] self.K_dis, self.L_dis = discretize( self.K_complex, self.L_complex, self.step ) def __call__(self, actions, start_state_rep): if len(actions.shape) == 2: return koopman_forward_single( self.K_dis, self.L_dis, self.action_encoder(actions), start_state_rep ) else: Vand_K = jnp.vander( self.K_dis, actions.shape[1] + 1, increasing=True ) return koopman_forward( Vand_K, self.L_dis, self.action_encoder(actions), start_state_rep )
@@
def causal_convolution(u, K): assert K.shape[0] == u.shape[0] ud = jnp.fft.fft(jnp.pad(u, (0, K.shape[0]))) Kd = jnp.fft.fft(jnp.pad(K, (0, u.shape[0]))) out = ud * Kd return jnp.fft.ifft(out)[: u.shape[0]] def compute_measurement_block(Ker, action_emb, init_state): return causal_convolution(action_emb, Ker[:-1]) + Ker[1:] * init_state def compute_measurement(Ker, action_emb, init_state): return jax.vmap( compute_measurement_block, in_axes=(0, 1, 0), out_axes=1 )(Ker, action_emb, init_state) def discretize(K, L, step): return jnp.exp(step*K), (jnp.exp(step*K)-1)/K * L def log_step_init(dt_min=0.001, dt_max=0.1): def init(key, shape): return jax.random.uniform(key, shape) * ( jnp.log(dt_max) - jnp.log(dt_min) ) + jnp.log(dt_min) return init def increasing_im_init(): def init(key, shape): return jnp.pi * jnp.arange(shape[0]) return init def random_im_init(): def init(key, shape): return jax.random.uniform(key, shape) return init @jit def koopman_forward(Vand_K, L, actions_emb, start_state_rep): action_emb_dim = actions_emb.shape[-1] state_dim = start_state_rep.shape[-1] actions_emb_complex = actions_emb[:, :, :action_emb_dim//2] \ + 1j * actions_emb[:, :, action_emb_dim//2:] start_state_rep_complex = start_state_rep[:, :state_dim // 2] \ + 1j * start_state_rep[:, state_dim // 2:] predicted_state_rep_complex = jax.vmap( compute_measurement, in_axes=(None, 0, 0), out_axes=0 )(Vand_K, actions_emb_complex @ L, start_state_rep_complex) predicted_state_rep = jnp.concatenate([ jnp.real(predicted_state_rep_complex), jnp.imag(predicted_state_rep_complex) ], -1) return predicted_state_rep @jit def koopman_forward_single(K, L, action_emb, start_state_rep): action_emb_dim = action_emb.shape[-1] state_dim = start_state_rep.shape[-1] action_emb_complex = action_emb[:, :action_emb_dim//2] \ + 1j * action_emb[:, action_emb_dim//2:] start_state_rep_complex = start_state_rep[:, :state_dim // 2] \ + 1j * start_state_rep[:, state_dim // 2:] predicted_state_rep_complex = K[None, :] * start_state_rep_complex + \ .action_emb_complex @ L predicted_state_rep = jnp.concatenate([ jnp.real(predicted_state_rep_complex), jnp.imag(predicted_state_rep_complex) ], -1) return predicted_state_rep